# Flowr.root – A flow matching based foundation model for joint multi-purpose structure-aware 3D ligand generation and affinity prediction

### v3

@@ Abstract
We present Flowr.root, an -equivariant flow-matching model for pocket-aware 3D ligand generation with joint binding affinity prediction and confidence estimation. The model supports multiple design modes including de novo generation, interaction/pharmacophore-conditional sampling, fragment elaboration, and multi-endpoint affinity prediction (pIC50, pKi, pKd, pEC50). Training combines large-scale ligand libraries with mixed-fidelity protein–ligand complexes, followed by refinement on curated co-crystal datasets and adaptation to project-specific data through parameter-efficient finetuning. Flowr.root achieves state-of-the-art performance in both unconditional 3D molecule and pocket-conditional ligand generation, producing geometrically realistic, low-strain structures with computational efficiency on established benchmark datasets. The integrated affinity prediction module demonstrates superior accuracy on the SPINDR test set and outperforms recent models on the Schrödinger FEP +/OpenFE benchmark while offering substantial speed advantages. As a foundation model, Flowr.root requires continuous parameter-efficient finetuning on project-specific datasets to account for unseen structure-activity landscapes, which we demonstrate yields strong correlation with experimental in-house data. The model’s joint generation and affinity prediction capabilities enable inference-time scaling through importance sampling, effectively steering molecular design toward higher-affinity compounds. Case studies validate this approach: selective CK2 ligand generation against CLK3 shows significant correlation between predicted and quantum-mechanical binding energies, while scaffold elaboration studies on ER, TYK2 and BACE1 demonstrate strong agreement between predicted affinities and QM calculations. By integrating structure-aware generation, affinity estimation, and property-guided sampling within a unified framework, Flowr.root provides a comprehensive foundation for structure-based drug design spanning hit identification through lead optimization.
1 Introduction
Diffusion- and flow-based generative models have emerged as leading frameworks in machine learning (ho2020ddpm; song2020sde; rombach2022ldm). Denoising Diffusion Probabilistic Models (DDPMs) learn a reverse process to transform noise into data samples (ho2020ddpm), while score-based models formulate generation via reverse-time stochastic differential equations (SDEs) or their probability-flow ordinary differential equation (ODE) counterparts (song2020sde). Flow matching, in turn, directly regresses the velocity field along a prescribed probability path to train continuous normalizing flows (lipman2022flowmatching; liu2022rectified), with stochastic-interpolant theory offering a unified perspective on diffusion and flow models (albergo2022building; albergo2023stochinterp). Beyond applications in natural language processing and computer vision, these models are increasingly applied to biology and chemistry, fueling advances in generative chemistry for drug discovery (model:edm; model:midi; model:eqgatdiff; model:diffsbdd; model:targetdiff; model:pilot; model:multiflow; model:semlaflow; model:flowr). In drug discovery, designing small molecules that selectively bind to specific protein targets remains a critical challenge. Diffusion and flow models have proven effective at capturing complex molecular and structural distributions, advancing methods from 3D molecular generation in structure-based drug design (SBDD) to protein–ligand (co-)folding (folding:alphafold3; boltz2_preprint).
One important subfield of AI-driven SBDD is pocket-conditional, structure-aware ligand design. Models like DiffSBDD (model:diffsbdd) and TargetDiff (model:targetdiff) have demonstrated that -equivariant diffusion models can generate ligands directly within protein pockets. Follow-up studies have focused on improving chemical validity and pose accuracy (model:eqgatdiff; model:molcraft) while coupling pocket conditioning with large-scale pretraining and multi-objective importance sampling to guide generation toward potency and synthesizability (model:pilot). Additionally, AI-driven models have increasingly integrated fragment priors into generative pipelines (imrie2020delinker; voloboev2024fragment2d; zhang2024fraggen; lee2025fragfm; guo2023linkinvent; model:difflinker). In fragment-based drug discovery (FBDD), which focuses on hit-to-lead optimization, models typically specialize either in de novo design or fragment-based strategies. However, practical lead optimization requires flexible navigation between these regimes, such as fragment growing or scaffold hopping under interaction constraints, which most current frameworks do not adequately address. Furthermore, many models do not incorporate the geometric constraints imposed by protein pockets. Recent work proposes a unified flow matching approach that integrates de novo, interaction-constrained, and fragment-based generation under pocket conditioning with efficient ODE sampling (model:flowr).
However, the effectiveness of generated ligands ultimately depends on their potency, which requires reliable affinity prediction. While experimental validation remains indispensable, computational prioritization during the design phase relies on accurate affinity predictions to effectively navigate target-relevant regions of chemical space. Classical scoring functions, such as AutoDock Vina (trott2010vina), Glide (friesner2004glide), and GOLD (jones1997gold), offer computational efficiency but often lack the necessary accuracy for reliable prioritization. Physics-based methods like free energy perturbation (FEP) and absolute binding free energy (ABFE) calculations provide higher precision (wang2015fepplus; mey2020bestpractices; mobley2012alchemical; alibay2022abfe; feng2022abfe; Ries2024_abfe), yet their computational cost prohibits application to large-scale generative campaigns, restricting their use to small subsets of candidates where absolute affinity data is critical. Machine learning–based scoring functions improve throughput but suffer from dataset bias and limited generalization (jimenez2018kdeep). Recent approaches using data augmentation have shown modest improvements (model:aev-plig), though they lack explicit structural information and remain insufficient for robust affinity prediction. The Boltz-2 ligand–protein co-folding model achieves near-FEP accuracy on selected targets with substantial speed advantages (boltz2_preprint). However, its decoupled architecture—where the affinity head is trained independently from the structural module—creates dependence on structure prediction quality, requires co-folding prior to each affinity prediction, and may limit co-adaptation between protein pocket and ligand geometry, particularly during project-specific fine-tuning.
These limitations motivate a framework that jointly learns structure and affinity while supporting fast, controllable, structure-aware generation. A model that learns the joint probability distribution over ligand geometry, binding pose, and affinity within the protein pocket context would enable efficient, potency-guided ligand generation with on-the-fly ranking capabilities. Crucially, joint training facilitates simple project-specific adaptation through fine-tuning—an essential capability given the distinct constraints of drug discovery campaigns, ranging from ADME/T requirements to R-group and scaffold novelty constraints, diverse assay readouts, and medicinal chemistry heuristics. A fundamental challenge in structure-based generative modeling lies in the inherent disconnect between public-domain training data and project-specific structure-activity relationships (SARs). While models can achieve broad coverage of chemical space through ligand generation, generalizing across novel bioactivity landscapes represents a fundamentally more complex problem. We posit that expecting universal generalization without adaptation is unrealistic; instead, models should function as dynamic companions that continuously refine their understanding of project-specific SARs through sustained interaction with incoming data. This paradigm shift necessitates moving from static models to efficient iterative refinement processes where model utility grows through continuous adaptation. Additionally, techniques such as multi-objective guidance via inference-time importance sampling (model:pilot) can help steering generation toward pre-specified, desired properties, while deeper distribution mismatches or activity cliffs can be addressed through direct preference alignment (drugflow) when suitable data is available.
While large ligand-only resources such as ZINC and PubChem offer abundant chemical diversity (dataset:zinc; dataset:pubchem), high-quality protein–ligand complexes with reliable affinity annotations remain scarce and noisy. Databases like PDBBind provide valuable training data (dataset:pdbbind), yet suffer from quality and coverage limitations; curated subsets such as HiQBind address these issues and even increase sample sizes (dataset:hiqbind), yet remain limited in scale. The Plinder dataset (dataset:plinder) expands scale by broadly curating the PDB, but most entries lack affinity annotations. More recent resources—BindingNet (dataset:bindingnetv2), Kinodata-3D (dataset:kinodata), and especially SAIR (dataset:sair)—improve chemical space coverage and include affinity data, though they do so with reduced accuracy. This heterogeneous data landscape however, motivates a multi-stage training paradigm: by systematically combining available resources ranked by fidelity, we can exploit their complementary strengths. Large-scale, lower-fidelity data establishes broad chemical space coverage and foundational structural understanding, while subsequent refinement on higher-quality, curated datasets sharpens affinity prediction and structural accuracy. Critically, this approach yields foundation models capable of efficient adaptation to project-specific objectives.
In this work, we present Flowr.root, a foundational framework for structure-based generative chemistry designed to address the continuum from large-scale pre-training through project-specific adaptation. Flowr.root integrates de novo, interaction- and pharmacophore-, as well as fragment-conditioned ligand generation with multi-endpoint affinity prediction and per-sample confidence estimation. Built on the FLOWR.multi backbone (model:flowr), the model employs a three-stage training strategy: (1) large-scale pre-training on billions of ligands and millions of protein–ligand complexes with diverse affinity labels, (2) fine-tuning on curated, higher-fidelity datasets, and (3) project-specific adaptation via parameter-efficient fine-tuning and multi-objective guidance through importance sampling. Critically, Flowr.root jointly trains its confidence and affinity prediction modules with structure generation, enabling affinity to actively shape the generative trajectory and facilitating rapid adaptation to project-specific data distributions. This design supports the model’s intended role as a continuously evolving companion throughout early-stage drug discovery campaigns—assisting in motif identification, core expansion, and R-group design through sustained refinement on incoming structure-activity data.
2 Related Works
Molecule generation
Early neural approaches to 3D molecular generation explored autoregressive models that sequentially add coordinates and atom types while maintaining geometric consistency. Symmetry-aware models such as G-SchNet (model:gschnet) and its conditional inverse-design follow-up (model:cgschnet) demonstrated that enforcing symmetries and conditioning signals can substantially improve the validity and controllability of generated 3D structures. Subsequent work proposed explicit autoregressive flows (model:autoregressive_flow). In parallel, the diffusion modeling paradigm matured from its non-equilibrium thermodynamics roots (diffusion:original) through score-based (song2020sde) and variational formulations (diffusion:ddpm; kingma2021vdm). These ideas rapidly translated to molecular geometry: xu2022geodiff and model:torsional_diffusion showed that -equivariant denoisers enable high-quality conformation generation by diffusing in Cartesian and torsional spaces, respectively.
The first -equivariant diffusion model for joint continuous coordinate and atom type generation was EDM (model:edm). Follow-ons pushed the design space along several axes: discrete and continuous combination with explicit bond order learning (model:midi), and enhanced denoising learning objectives (model:eqgatdiff), establishing strong baselines.
Thereafter, diffusion-based methods pioneered pocket-aware de novo design, also utilizing -equivariant networks. Notably, DiffSBDD (model:diffsbdd) and TargetDiff (model:targetdiff) demonstrated that conditional diffusion models can generate diverse, target-specific ligands within the protein pocket, adhering to symmetry constraints and enabling task versatility through sampling controls. More recently, PILOT (model:pilot) combined large-scale pretraining, pocket conditioning, and property guidance, highlighting the importance of multi-objective steering (e.g., drug-likeness, synthesizability) under structure constraints. Together, these works established that pocket-aware 3D diffusion can simultaneously respect symmetry, improve pose realism, and support versatile constraints via conditioning and guided sampling.
Meanwhile, flow matching (lipman2022flowmatching) (FM) was proposed refining continuous normalizing flows by directly regressing a time-dependent velocity field that pushes a simple prior to the data distribution, offering faster sampling and flexible priors. SemlaFlow (model:semlaflow) introduced a scalable -equivariant architecture (Semla) trained via flow matching, achieving state-of-the-art unconditional 3D molecule generation with significant speed-ups. Pushing FM into SBDD, FLOWR (model:flowr) extended the Semla-style backbones with an dedicated pocket encoder and mixed continuous/categorical FM, supporting multi-mode de novo, interaction-guided, and fragment-based generation in a single model. FLOWR reports large speedups over pocket-diffusion baselines.
Fragment-based drug discovery (FBDD) motivates models that initiate from fragments and perform growth, linking, or merging under pocket constraints. Early deep generative linker design incorporated 3D information into graph models (DeLinker) imrie2020delinker, while SyntaLinker and AutoLinker utilized conditional transformers to synthesize linkers directly in the SMILES space, given fragment pairs and constraints yang2020syntalink; feng2022syntalinkhybrid. Recent advances have introduced -equivariant models: DiffLinker formulates linker generation as an -equivariant conditional diffusion, explicitly learning 3D geometry between fragment anchors model:difflinker. For broader medicinal-chemistry workflows, Link-INVENT extends REINVENT with reinforcement learning to optimize linkers for multiple objectives, demonstrated on fragment linking, scaffold hopping, and PROTAC design guo2023linkinvent. STRIFE extracts target-specific pharmacophoric features to steer elaboration in 3D hadfield2022strife, while AutoFragDiff integrates fragment-wise, autoregressive diffusion with pocket conditioning to improve local 3D geometry during growth ghorbani2023autofragdiff. For scaffold hopping, DiffHopp employs an -equivariant graph diffusion model tailored for scaffold replacement conditioned on a protein–ligand complex torge2023diffhopp, and TurboHopp accelerates pocket-conditioned 3D scaffold hopping with consistency models and reinforcement learning-based preference optimization yoo2024turbohopp.
Binding Affinity Prediction
Estimating the change in free energy upon binding (, or affinity) accurately remains a cornerstone of structure-enabled small-molecule discovery. Binding affinity is relevant for all early stages of drug discovery, starting from hit identification, where the goal is to find tight and selective binders, through hit-to-lead and lead optimization, where potency must be balanced with absorption, distribution, metabolism, excretion, safety, toxicity, and efficacy considerations. Given the astronomical size of chemical space, computer-aided drug design (CADD) is indispensable to select and prioritize candidates in silico before spending scarce experimental resources (reymond2010chemicalspace; jorgensen2004manyroles). Classical structure-based approaches to affinity prediction span knowledge-based scoring and physics-based models grounded in molecular mechanics (liu2015classification; gilson2007calculation). Heuristic docking scores offer speed at the expense of physical rigor. Empirical scoring functions such as AutoDock Vina, Glide, or GOLD remain widely used due to speed, but show inconsistent results between targets trott2010vina; friesner2004glide; jones1997gold. Semi-empirical and QM / MM scoring have closed part of the gap at an intermediate cost, for example, SQM2.20 achieves DFT-quality affinity estimates in minutes, but only on selected targets (pecina2024sqm220; molani2024qmmm). End-point methods such as MM-PBSA and MM-GBSA combine molecular mechanics with continuum solvation to approximate the from MD snapshots at a relatively low cost and remain widely used when throughput is critical (kollman2000accresearch; homeyer2012mmpbsa; still1990gb; gohlke2004rasraf). Alchemical binding free energy methods, absolute (ABFE) and relative (RBFE), trade throughput for accuracy (aldeghi2016_abfe; boresch2003abfe; gilson1997statmech; Fu2022_abfe; feng2022abfe; cournia2017rbfe; Ross2023_rbfe). Modern workflows based on the free energy perturbation theory (FEP) (zwanzig1954perturbation) have achieved impressive accuracy on suitable congeneric series (abel2017enhanced; wang2015fepplus; ross2023maximal), but remain sensitive to force fields and system preparation.
[…]

### v5

@@ Abstract
We present Flowr.root, an -equivariant flow-matching model for pocket-aware 3D ligand generation with joint potency and binding affinity prediction and confidence estimation. The model supports de novo generation, interaction- and pharmacophore-conditional sampling, fragment elaboration and replacement, and multi-endpoint affinity prediction (pIC50, pKi, pKd, pEC50). Training combines large-scale ligand libraries with mixed-fidelity protein–ligand complexes, refined on curated co-crystal datasets and adapted to project-specific data through parameter-efficient finetuning. The base Flowr.root model achieves leading performance in unconditional 3D molecule and pocket-conditional ligand generation, producing geometrically realistic, low-strain structures with high efficiency. On HiQBind, the pre-trained and finetuned model demonstrates highly accurate affinity predictions, and outperforms recent state-of-the-art methods such as Boltz-2 on the FEP+/OpenFE benchmark with substantial speed advantages. However, we show that addressing unseen structure–activity landscapes requires domain adaptation; parameter-efficient LoRA finetuning yields marked improvements on diverse proprietary datasets and PDE10A. Joint generation and affinity prediction enable inference-time scaling through importance sampling, steering design toward higher-affinity compounds. Case studies validate this: selective CK2 ligand generation against CLK3 shows notable correlation between predicted and quantum-mechanical binding energies. Scaffold elaboration on ER, TYK2, and BACE1 demonstrates strong agreement between predicted affinities and QM calculations while confirming geometric fidelity. By integrating structure-aware generation, affinity estimation, property-guided sampling, and efficient domain adaptation, Flowr.root provides a comprehensive foundation for structure-based drug design from hit identification through lead optimization.
Introduction
Diffusion- and flow-based generative models have become central tools for learning complex data distributions and have recently enabled rapid progress in generative chemistry and structure-based drug design (SBDD) [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 4, 44, 42, 43, 17]. One key subfield is pocket-conditional ligand generation, where -equivariant models generate ligands directly in the protein binding site [40, 41]. Subsequent work has improved chemical validity and pose quality and explored pre-training and inference-time guidance to steer sampling toward desired objectives [39, 49, 4]. Complementary to de novo design, fragment-based approaches support practical lead optimization; however, combining pocket conditioning with flexible fragment elaboration and editing (for example, scaffold replacement or fragment growing under interaction constraints) within a single generative backbone remains challenging [50, 51, 52, 53, 54, 55, 17].
Beyond generation, the utility of proposed ligands depends on reliable potency and binding affinity prediction to prioritize candidates during iterative design. Classical scoring functions (for example, docking) are computationally efficient but can be insufficiently accurate or are not representing affinities, whereas more elaborate physics-based methods such as free energy perturbation (FEP) and absolute binding free energies (ABFE) are typically more reliable but too costly to apply at the scale of modern generative campaigns [56, 57, 58, 59, 60, 61, 62, 63, 64]. Machine learning scoring functions improve throughput but may suffer from dataset bias and limited generalization [65, 111, 24]. Co-folding approaches such as Boltz-2 can reach high accuracy on selected targets, but require structure prediction per query and, due to decoupled training, may limit co-adaptation between pocket geometry and affinity [45].
These considerations motivate models that jointly learn (conditional) structure generation and affinity prediction in a protein pocket context. Joint modeling supports in-distribution ranking of generated structures and enables inference-time steering (for example, via importance sampling) without relying on external scoring functions [4]. At the same time, drug discovery campaigns are characterized by assay- and series-specific structure–activity relationships (SARs) that often differ from public training data. Consequently, strong benchmark performance does not necessarily translate to a new project without adaptation; efficient finetuning and, where appropriate, preference-based alignment [66], can help align the model to the relevant local SAR landscape.
Training such models is further constrained by the scarcity of high-quality protein–ligand complexes with reliable affinity annotations relative to ligand-only resources. This motivates multi-stage training strategies that exploit large-scale, lower-fidelity data to learn broad chemical and structural priors, followed by refinement on curated experimental datasets to improve structural accuracy and affinity prediction [2, 3, 67, 19, 1, 11, 9, 12].
In this work, we present Flowr.root, a foundation model for structure-based drug design that spans large-scale pre-training to project-specific adaptation. Within a single architecture, Flowr.root unifies de novo, interaction- and pharmacophore-conditional, and fragment-conditioned ligand generation with multi-endpoint affinity prediction and confidence estimation. The model employs a three-stage training strategy that leverages datasets of varying fidelity: (1) large-scale pre-training on billions of ligands and millions of protein–ligand complexes with diverse affinity labels, (2) refinement on curated, higher-quality structural data, and (3) project-specific adaptation via parameter-efficient Low-Rank Adaptation (LoRA) [47] and multi-objective guidance through importance sampling [4]. This joint training paradigm allows affinity prediction to directly guide the generative process without external scoring functions, while supporting rapid adaptation to project-specific data. In addition, we introduce flexible fragment-conditional generation modes—including local fragment growing and replacement for spatially targeted modifications enabled by a mixed isotropic-anisotropic prior placement strategy. Combining joint structure and affinity prediction with fragment-based elaboration, affinity guidance enables effective sampling and ranking of local chemical spaces. On established benchmarks, Flowr.root achieves PoseBusters-validity for pocket-conditional ligand generation and Pearson correlations of up to for binding affinity prediction on the FEP+/OpenFE benchmark. Together, these capabilities position Flowr.root as a versatile tool for early-stage drug discovery, from hit identification through lead optimization.
Results
Overview of Flowr.root
Flowr.root is a unified framework for structure-based drug design that jointly trains pocket-aware 3D ligand generation with multi-endpoint affinity prediction and confidence estimation (Fig. 1). The framework addresses key challenges in computational drug discovery: generating chemically valid, geometrically realistic ligands within protein binding sites; accurately predicting potency and binding affinity across diverse assay types; supporting flexible generation modes from de novo design to targeted fragment modifications; and enabling efficient adaptation to project-specific structure-activity relationships. A comprehensive overview is given in Methods.
The model is built on an -equivariant flow matching backbone that learns a mixed continuous/discrete transport map, transforming a prior distribution (noise or fragment anchors for coordinates, uniform for categorical features) to the target ligand distribution within a given protein pocket. The architecture comprises two main components: a pocket encoder that processes full-atom protein features through equivariant self-attention layers, generating invariant and equivariant representations; and a ligand decoder that employs equivariant self-attention for intra-ligand dependencies followed by cross-attention to integrate pocket context. Protein pockets are extracted using a 7Å cutoff radius around binding sites, balancing computational tractability with interaction coverage. The ligand decoder features three output heads that serve complementary purposes. The structure head predicts atomic coordinates, atom types, bond orders, partial charges, and hybridization states. The multi-affinity head comprises four separate predictors for pIC50, pKi, pKd, and pEC50—explicitly avoiding treatment of heterogeneous affinity labels as interchangeable. The confidence head provides pLDDT-based uncertainty estimation for generated structures.
Training follows a three-stage paradigm that exploits data of varying fidelity to address the scarcity of high-quality experimental complexes: Stage 1 builds broad generative priors across chemical and structural distributions. The model trains on 1.5B small molecule conformations (ZINC3D, PubChem3D, Enamine REAL, OMol25) for chemical coverage, and 2.5M mixed-fidelity protein-ligand complexes with and without affinity labels from both computational sources (BindingNet, SAIR, KIBA-3D, Davis-3D, Kinodata-3D) and experimental sources (Plinder, BindingMOAD). Stage 2 adapts pre-trained weights to curated, drug-like datasets. We use SPINDR and HiQBind for multi-target finetuning, teaching the model fine details of protein-ligand interactions and binding affinity from high-quality co-crystal structures. Stage 3 bridges the gap between public data and project-specific structure-activity landscapes through multiple strategies: inference-time scaling via importance sampling guides generation toward desired properties (affinity, ADME/T, synthetic accessibility); parameter-efficient finetuning enables rapid adaptation while avoiding catastrophic forgetting. This multi-stage paradigm addresses a fundamental limitation: strong benchmark performance does not guarantee generalization to unseen structure-activity landscapes. Real-world drug discovery requires understanding nuanced SAR across related molecules within specific binding sites—richness uniquely captured by dense bioactivity data that differs substantially from sparse public datasets. Rather than expecting universal generalization, Flowr.root is designed as an adaptive companion that refines through sustained interaction with project-specific data.
The model is trained using mean squared error (MSE) loss for coordinates and categorical cross-entropy (CE) losses for atom types, charges, hybridization, and bond types. In addition, we introduce explicit geometric supervision through bond length and bond angle losses. The bond length loss penalizes deviations in predicted inter-atomic distances for bonded pairs. The bond angle loss uses Huber loss over valid angle triplets to preserve local molecular geometry while remaining robust to outliers. These auxiliary losses substantially reduce strain energies in generated structures. For training the affinity head, invariant and equivariant features from ligand, pocket, and their interactions via the ligand decoder’s latent representations are extracted, and gated equivariant blocks combine invariant and equivariant tensors, followed by multi-layer perceptrons (MLPs) that produce aggregated ligand, pocket, and interaction embeddings. Task-specific heads then predict each affinity type separately, with Huber loss for robust training. This design explicitly models the distinct experimental setups underlying IC50, Ki, Kd, and EC50 measurements rather than treating them as interchangeable.
Flowr.root supports multiple generation modes within a single backbone, enabling flexible application across hit identification and lead optimization: (1) de novo pocket-conditional generation for exploring diverse chemical matter; (2) interaction/pharmacophore-conditional generation for preserving or enforcing specific protein-ligand contacts; (3) scaffold hopping and elaboration for core and R-group replacement; and (4) fragment growing and local fragment replacement for targeted structural modifications. For the latter mode, we use a flexible prior placement strategy that shifts the generative prior from the zero center-of-mass to the local replacement site, enforcing locality in partial modifications and enabling precise control over which atoms to replace while preserving the remainder of the molecule. In addition, we propose a mixed isotropic and anisotropic Gaussian prior placement during training to enable a more targeted ligand generation strategy at inference. We provide more details in Methods. Flowr.root also supports inference-time steering via importance sampling to sample from conditional distributions allowing for single- and multi-objective optimization without retraining, for example, steering toward higher predicted affinities while maintaining sample diversity.
To train Flowr.root, we leveraged a diverse collection of datasets. A detailed description of all datasets used in this work, including preprocessing and curation pipelines, dataset statistics, and chemical space analyses, is provided in Methods. In short, all protein-ligand complexes undergo rigorous preprocessing using Schrödinger’s LigPrep and PrepWizard, including protonation state determination at physiological pH, protein preparation with side chain completion, and constrained minimization. Protein pockets are extracted using a 7Å cutoff radius around reference ligands. Crucially, if not stated otherwise, throughout all protein-ligand datasets we kept a consistent dataset split following the provided Plinder [1] train, validation and test set splits to avoid data leakage. Plinder employs comprehensive similarity metrics (protein sequence, pocket-level Jaccard, interaction-level PLIP, and ligand-level Tanimoto) to ensure minimal leakage and high-quality, non-redundant testing. To our knowledge, this represents the most rigorous publicly available splitting strategy for structure-based modeling.
Unconditional and Pocket-Conditional Generation
We first establish the improvements conferred by the architectural changes incorporated into Flowr.root. To this end, we demonstrate the expressiveness of the ligand decoder backbone for unconditional 3D molecule generation on the well-established Geom-Drugs dataset, comparing against recent state-of-the-art models including Eqgat-diff [39], SemlaFlow [43], ADiT [48], Megalodon [49], and FlowMol3 [48]. Here, our non-pretrained base model, Flowr.root, achieves 0.94 PoseBusters-validity (mean), surpassing all baselines including the recently released FlowMol3 model (0.92), with a median relaxation energy of 3.65 kcal/mol and a median relaxation RMSD of only 0.07 Å—demonstrating high geometric precision. Complete benchmark results and comparisons are provided in Supplementary Table 1.
For pocket-conditional ligand generation, we benchmark on CrossDocked2020, where Flowr.root achieves a mean PoseBusters-validity of and a mean strain energy of kcal/mol, both of which are substantially lower than all competing models, including Flowr (mean PoseBusters-validity of and mean strain energy of kcal/mol) as well as the next best model Pilot (mean PoseBusters-validity of and strain energy of kcal/mol), and the highest mean AutoDock-Vina score of kcal/mol (vs. and kcal/mol). Complete benchmark results and comparisons are provided in Supplementary Table 2.
Pocket-Conditional Ligand Generation: Spindr
Model PB-valid Strain energy Vina score Vina score BondA.W1 BondL.W1 [] DihedralW1 Pilot 0.79 .21 120.10 71.61 -6.30 .96 -6.68 1.07 1.82 0.42 5.52 Flowr 0.93 .22 90.05 52.18 -6.93 .92 -7.22 .92 1.08 0.35 3.88 Flowr.root 0.97 .10 50.36 34.59 -7.52 .84 -7.71 .85 0.60 0.43 3.62 Test set 0.99 .04 43.27 41.85 -7.69 2.00 -7.88 2.00 - - -
Next, we evaluate Flowr.root’s ability to generate valid, physically plausible ligands, but on the more challenging test set of the Spindr dataset. We compare the non-pretrained Flowr.root model with Flowr and the diffusion-based Pilot model. Table 1 reveals that Flowr.root substantially outperforms both Flowr and Pilot across all metrics. Flowr.root achieves a mean PoseBusters-validity of , nearing the test set reference (). Additionally, the strain energy statistics substantially decreases to kcal/mol, further demonstrating the model’s ability to generate chemically valid and energetically favorable ligands. Flowr.root also performs better in docking accuracy, with a mean Vina score of kcal/mol.
In addition, we analyzed a wide range of complementary metrics (Supplementary Table 3-6). Flowr.root achieves 100% novelty () with high diversity () and uniqueness () rates, comparable to Flowr, while demonstrating consistent and substantial performance improvements across fragment-conditional generation modes, including scaffold hopping and pharmacophore-conditional generation.
Affinity prediction: HiQBind
A central part of Flowr.root is its joint structure–affinity modeling. In Fig. 2 we provide a detailed analysis of affinity prediction performance on the HiQBind test set. The affinity head of Flowr.root provides accurate predictions across various affinity types, including pIC50, pKi, and pKd. For pIC50, the model achieves a Pearson correlation of and an R2 of , while for pKi and pKd we get Pearson correlations of and , respectively. When aggregating predictions across all affinity types (median), the model achieves RAE , RMSE , R2 , Kendall , and Pearson . Note, reported values are mean values across four seed runs the mean 95% confidence interval (CI) across seeds. These results confirm that the affinity head, trained jointly with structure generation, provides accurate and reliable potency estimates with robust performance across different experimental affinity labels. Probing the Flowr.root model trained solely on HiQBind without pre-training, we observe substantially worse performance across all metrics with RAE , RMSE , R2 , Kendall , and Pearson underlining the importance of the proposed large-scale pre-training pipeline. Note, in both cases we strictly follow the Plinder dataset split minimizing information leakage between train and test also when pre-training.
Beyond predictive accuracy, Flowr.root enables property-guided generation via inference-time steering. Fig. 3 illustrates the effect of importance sampling-based steering on the predicted affinity distribution of generated ligands on the HiQBind test set. Without guidance, the mean predicted pIC50 is . As the steering duration increases (from to of the trajectory), the mean predicted pIC50 shifts progressively to higher values (, , ), while the standard deviation remains stable. This demonstrates that the model can be effectively biased toward predicted higher-affinity ligands during generation.
In Fig. 3, we show the distribution of important chemical properties comparing unguided with guided ( steering duration) samples across targets on the HiQBind test set. While the distribution of SA scores shows a slight decrease towards lower scores for guided samples, we find that the overall druglikeness increases substantially with a mean Lipinski score of 4.30 for the unguided to 4.68 for the guided samples. Interestingly, we find that guidance consolidates the sample space, leading to mostly more compact distributions. This can also be seen in the chemical space comparison in Fig. 3b–c, where we overlay the chemical spaces of the unguided and guided samples using both PCA and UMAP on molecular fingerprints.
Lastly, we also analyze geometric validity and sample diversity as steering intensity increases. For unguided versus guided samples, we observe a mean PoseBusters-validity of 0.93 versus 0.91, mean strain energy of 48.46 versus 54.42 kcal/mol, mean Vina score of -7.41 vs. -7.50, and mean per-target diversity of 0.79 versus 0.79. While we observe a slight decrease in PoseBusters-validity and increase in strain energies, the diversity of the samples remains the same, so we can rule out a potential mode collapse when using guidance. In addition, we see a slight decrease in Vina scores on guided samples, demonstrating that we can maintain physically realistic structures while optimizing for affinity.
Affinity prediction: Schrodinger FEP+ and OpenFE
Next, we evaluate the performance of the affinity prediction module of Flowr.root using the Schrodinger FEP+/OpenFE IndustryBenchmark datasets [51], comparing Flowr.root to state-of-the-art methods. These benchmarks offer a robust test of the model’s prediction capabilities with experimentally determined binding affinities across various targets and affinity types for protein-ligand complexes.
In Fig. 4a, we present the scatter plot for Flowr.root-predicted combined affinity predictions (median over pKd, pKi, and pIC50) versus experimental binding affinities. The plot indicates strong correlation with minimal deviation, suggesting that Flowr.root can successfully predict binding affinities across a broad range of protein-ligand complexes. In Fig. 4b, we show the performance of Flowr.root using various error and correlation metrics in comparison to OpenFE, FEP+, AEV-PLIG and Boltz-2. Flowr.root achieves an RMSE of kcal/mol, with a Pearson correlation of and a Kendall of . These results surpass those of FEP+ (RMSE , Pearson ) and OpenFE (RMSE , Pearson ), as well as those of Boltz-2 (Kendall , Pearson ) and AEV-PLIG (RMSE , Kendall ). Note, reported values for Flowr.root are mean values across four seed runs the mean 95% CI. For FEP+ and OpenFE, the mean 95% CI is shown [52], while Boltz-2 and AEV-PLIG do not provide confidence intervals.
Flowr.root outperforms physics-based models, namely OpenFE and FEP+, on almost all error and correlation metrics and substantially outperforms recent state-of-the-art ML-based models AEV-PLIG and Boltz-2 across all metrics. Notably, Flowr.root achieves this while being 3x faster than AEV-PLIG, 200x faster than Boltz-2, and over 10000x faster than OpenFE and FEP+. However, we also want to emphasize that the Schrodinger FEP+/OpenFE dataset is likely substantially overlapping with the training data in terms of ligand and target space. While this applies to AEV-PLIG and Boltz-2 as well allowing for a somewhat fair comparison, we do not expect these results to indicate substantial generalization capabilities.
Domain Adaptation via Finetuning: In-House Data
Next, we investigate the performance of Flowr.root on proprietary project-specific data. We evaluate four diverse pharmaceutical projects spanning inflammatory and neurodegenerative disease, to oncology, each encompassing distinct protein targets and ligand chemical spaces ranging from conventional small molecules to molecular degraders. Dataset sizes vary across projects: Project 1 comprises 1,073 compounds, Project 2 contains 730 compounds, Project 3 includes 229 compounds, and Project 4 contains 618 compounds. Projects 1–3 are annotated with pIC50 values, while Project 4 uses pEC50 measurements.
Pharmaceutical projects typically generate substantially more bioactivity data than co-crystal structures. To leverage existing assay data, we developed a structure generation workflow using OpenEye’s docking tools [53]. Measured compounds were first matched to existing co-crystal ligands via maximum common substructure (MCS) alignment and assigned to protein structures with the highest MCS similarity. Depending on the TanimotoCombo similarity between the query molecule and reference ligand (threshold TC , computed via ROCS overlays [54] on up to 800 Omega-generated conformers [55]), we employed either Hybrid or ShapeFit docking. Docked poses were subsequently minimized using OpenEye’s Szybki with the ff14sb force field [56, 57], and poses with positive interaction energies were discarded. The top pose was selected based on RMSD to the reference ligand.
We evaluate both Boltz-2 and Flowr.root in zero-shot settings on held-out test sets comprising 100 randomly selected complexes per project (80 for Project 3 due to its smaller size), with the remaining samples used for training and validation. As shown in Fig. 5a–b, both models fail to generalize to these unseen bioactivity spaces, yielding negligible or negative correlations between predicted and experimental potencies. Specifically, Boltz-2 yields strongly negative across all projects; for Flowr.root, values are (Project 1), (Project 2), (Project 3), and (Project 4). These results highlight a fundamental limitation: zero-shot generalization across arbitrary structure–activity landscapes is unrealistic given the inherent complexity of structure–activity relationships.
[…]